How quickly can you do this? Fill appropriate sign. ( ‘’) 
3 / 4 ☐ 7 / 8

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

3 / 4 = (3 × 2) / (4 × 2)
= 6 / 8
Between 6 / 8 and 7 / 8
Both have same denominators. So, the fraction having greater numerator will be greater
∴3 / 4 < 7 / 8

Thanks